OpenStudy (anonymous):

I have the answer can someone just check me please.... Find the derivative of the function h(x)=(4x^2-e^x)^8

OpenStudy (amistre64):

show me your work :)

OpenStudy (anonymous):

ok I got h(x)=8(4x^2-e^x)^7(8x-e^x)

OpenStudy (amistre64):

loks good to me h(x)=(4x^2-e^x)^8 h'(x) = 8(4x^2-e^x)^7 (4x^2-e^x)' h'(x) = 8(4x^2-e^x)^7 (8x-e^x)

OpenStudy (anonymous):

awesome thanks

OpenStudy (amistre64):

youre welcome :)